# LagCeresa.bib

### v1

@@ 1. Introduction
[zharkov_tropical_ceresa, Theorem 3] For a generic genus $3$ tropical curve of type $K4$, $C$ and $C^{-}$ are not algebraically equivalent.
A tropical curve of type $K4$ is represented in Figure 2, and is only one of five types of generic genus $3$ tropical curves. The other four are drawn in Figure 7. It distinguishes itself as the only non-hyperelliptic type; this is explained and discussed in Section 8.1. The fact that Ceresa’s classical statement does not translate for other generic types can be understood as a consequence of the failure of the tropical Torelli theorem.
Zharkov’s proof of Theorem 1.2 introduces a specific instance of what we refer to in this paper as tropical flux, namely a morphism
$\Theta:\text{Griff}_{1}^{trop}(J(C))\longrightarrow\mathbb{R}/\mathcal{P},$
where $\mathcal{P}$ is a finitely generated subgroup in $\mathbb{R}$. This morphism is given by integrating a tropical differential form $\Omega_{0}$, called the determinantal form, over a $2$-chain bounding cycles in $\text{Griff}_{1}^{trop}(J(C))$. The lattice $\mathcal{P}$ is simply given by the periods of $\Omega_{0}$. One can then conclude by proving that, for generic $C$,
$m\cdot\Theta(C-C^{-})\neq 0\mod\mathcal{P}$
for any $m\in\mathbb{Z}$. In Section 5.3 we extend this to aribtrary $k$-cycles in a tropical torus $B$ by defining morphisms
$\Theta_{k}:\text{Griff}_{k}^{trop}(B)\longrightarrow\mathbb{R}/\mathcal{P}_{k}$
which we call tropical fluxes.
Our proof of Theorem 1.3 is based on the observation that, in the Lagrangian torus fibration
$X(J(C)):=T^{*}J(C)/T^{*}_{\mathbb{Z}}J(C)\longrightarrow J(C)$
associated to the tropical $3$-torus $J(C)$, the tropical flux $\Theta$ is related to the morphism
$\mathcal{C}ob^{or}(X(J(C)))_{hom}\longrightarrow\mathbb{R}/\left([\omega^{2}]\cdot H_{4}(X(J(C));\mathbb{Z})\right)$
$\hskip 31.29802ptL\hskip 36.98857pt\mapsto\hskip 45.5244pt\int_{\gamma}\omega^{2},$
where $\mathcal{C}ob^{or}(X)_{hom}$ denotes the homologically trivial subgroup of the oriented cobordism group $\mathcal{C}ob^{or}(X)$, and $\gamma$ is a $4$-chain whose boundary is $\partial\gamma=L$. The content of Section 2 is to exhibit this as a specific instance of symplectic fluxes, which are morphisms
$\mathcal{C}ob^{or}(X^{2n})_{hom}\longrightarrow\mathbb{R}/\mathcal{Q}$
built from certain differential forms over $(n+1)$-chains bounding $L\in\mathcal{C}ob^{or}(X^{2n})_{hom}$. Here $\mathcal{Q}$ is the period lattice of the corresponding differential form.
The proof of Theorem 1.3 can now be sketched as follows. Still taking $C$ to be a genus $3$ tropical curve of type $K4$, we define the Lagrangian Ceresa cycle $L-L^{-}\in\mathcal{C}ob^{or}(X(J(C)))_{hom}$ as a Lagrangian lift of the tropical Ceresa cycle $C-C^{-}\in\text{Griff}_{1}^{trop}(J(C))$ see 7.1 for the description of $L$ and $L^{-}$ as $3$-manifolds. We then show that $\mathcal{P}$ coincides with the periods of $\omega^{2}$, and that
$\int_{\gamma_{0}}\omega^{2}=\Theta(C-C^{-})\in\mathbb{R}/\mathcal{P}$
for some $\gamma_{0}$ whose boundary is $\partial\gamma_{0}=L-L^{-}$, allowing us to conclude.
In other words, we exhibit an instance of the following commutative diagram:
@@
therefore the Ceresa cycle in $\text{Griff}_{1}(J(C))_{hom}$ having non-torsion image implies that its Lagrangian lift does as well.
Structure of the paper
In Section 2 we give the definition of the various flavours of Lagrangian cobordisms. We elaborate on the notion of symplectic fluxes and their relation to (oriented) Lagrangian cobordism groups as obstructions to nulhomologous Lagrangians being nullcobordant. We include an example application to the cotangent bundle of a torus.
Section 3 introduces and motivates the notion of algebraic Lagrangian cobordism, illustrated by simple examples. We then show that certain symplectic fluxes are well-defined from algebraic Lagrangian cobordism groups, and that these fluxes can be encoded in a map to a Lefschetz Jacobian, a naturally tropical object.
Section 4 collects relevant tropical background; it contains short expositions on tropical tori, tropical curves, their Jacobians and Ceresa cycles, as well as tropical (co)homology for tropical tori.
In Section 5, we introduce the notion of tropical fluxes for tori, which are defined using tropical (co)homology. These are given by integrals of differential forms, which are generalisations of Zharkov’s determinantal form. They detect obstructions to a nullhomologous tropical cycle being tropically algebraically trivial.
The aim of Section 6 is to provide the technical tools allowing one to translate tropical fluxes into symplectic fluxes in the associated Lagrangian torus fibration. Because both types of fluxes are given by integration of certain differential forms, this involves constructing maps of (co)chains from tropical (co)chains in $B$ to singular (co)chains in $X(B)$ in a way that preserves the integration pairing. The underlying maps on (co)homology are a Künneth-type decomposition for $X(B)\cong B\times T^{n}$.
We prove our main Theorem 1.3 in Section 7, after an explicit construction of the Lagrangian Ceresa cycle as a $3$-manifold inside a symplectic $6$-torus. We speculate on a higher-dimensional extension of this result.
We end with two additional remarks in Section 8. The first turns to the case of genus $3$ tropical curves of hyperelliptic type, whose Ceresa cycles and corresponding tropical fluxes provide one with a rich source of interesting questions and have inspired recent work [ceresa_period_tropical, Ceresa_class_corey_et_al]. The second relates our work to a construction by Reznikov [Reznikov_characteristic_classes_in_symplectic_topology] building characters of symplectic Torelli-type groups.
Notation
We abuse notation and simply denote by $X$ a symplectic manifold $(X,\omega)$. Similarly, $(X,-\omega)$ will be written $\overline{X}$.
Acknowledgements
I am greatly endebted to Ivan Smith for his guidance and support throughout this project, and for his careful reading of previous drafts of this paper. I was supported by EPSRC grant EP/X030660/1 for the duration of this work.

### v2

@@ 2.1.3. An important example: the genus $3$ curve of type $K4$
The fact that Zharkov’s argument does not apply to other non-hyperelliptic curves which are of hyperelliptic type, can be understood as a failure of the tropical Torelli Theorem; this is discussed in Section 7.1.
Remark 2.6.
Theorem 2.5 is only non-trivial because the Ceresa cycle is nullhomologous. In the classical case, this follows from the fact that (1)(-1) acts trivially on $H_{j}(J(C);\mathbb{Z})$ for $j$ even. Here, it follows immediately from the fact that Zharkov constructs an explicit $2$-chain $\gamma_{0}$ with $\partial\gamma_{0}=C-C^{-}$.
Remark 2.7.
The stronger statement that $C-C^{-}$ has infinite order in $CH_{1}(J(C))$ modulo algebraic equivalence follows from Zharkov’s proof, and is proved by Ceresa in the classical setting [Ceresa]. We address this again in Remark 6.4.
Let us now study this curve and its Jacobian more explicitly. The basis $\{\gamma_{1},\gamma_{2},\gamma_{3}\}$ of $H_{1}(C;\mathbb{Z})=\Gamma_{1}$ is represented in Figure 3. Recall that $\Gamma_{2}\subset H^{0}(T^{*}_{\mathbb{R}}C)^{*}\cong\mathbb{R}^{3}$ is the lattice of those maps which evaluate to integers on elements of $H^{0}(T_{\mathbb{Z}}^{*}C)$. This determines the tropical structure of $J(C)$. A basis $\{e_{1},e_{2},e_{3}\}$ of $\Gamma_{2}$ is also represented by the vectors in Figure 3. These two bases are related through
Figure 3. Bases for $\Gamma_{1}$ and $\Gamma_{2}$ represented in $C$.
$a=1$ $b=0.5$ $c=1.5$ $d=1$ $e=0.5$ $f=1.$
@@
Figure 4. A representation of the embedding of the type $K4$ tropical curve in its Jacobian, with the edges colour-coded as in Figure 6, and the two bases $\{\gamma_{1},\gamma_{2},\gamma_{3}\}$ and $\{e_{1},e_{2},e_{3}\}$ represented as well. The two black parallelepipeds are two copies of $J(C)$ inside $\Gamma_{2}\otimes\mathbb{R}\cong\mathbb{R}^{3}$, and edges with the same symbols $\blacksquare,\blacktriangle$, or $\bigstar$ are identified.
2.2. Tropical flux
Recall from the Introduction the main ideas of Zharkov’s proof of Theorem 2.5. He starts by introducing a $2$-form $\Omega_{0}$ on $J(C)$ called determinantal form. He then shows that, if two $1$-cycles are algebraically equivalent (in particular they are homologically equivalent), then integrating $\Omega_{0}$ over any $2$-chain realising this homological equivalence must yield a period of $\Omega_{0}$. He proceeds to constructing an explicit $2$-chain $\gamma_{0}$ in $J(C)$ with $\partial\gamma_{0}=C-C^{-}$, computes $\int_{\gamma_{0}}\Omega_{0}$, and shows it is not a period of $\Omega_{0}$ for generic $C$. The aim of the current Section is to formalise and generalise this idea of using non-trivial integrals to detect obstructions to algebraic equivalence. The result are morphisms
$\Theta_{k}:\text{Griff}_{k}(B)\longrightarrow\mathbb{R}/\mathcal{P}_{k},$
where $\text{Griff}_{k}(B)$ is the $k$-th Griffiths group of $B$ of homologically trivial $k$-cycles modulo algebraic equivalence, and $\mathcal{P}_{k}$ is a corresponding set of periods. This involves generalising the determinantal form $\Omega_{0}$, and the construction of certain $(k+1)$-chains bounding tropical $k$-cycles. The objects involved are tropical chains and cochains; their coefficients encode information about the tropical structure of the underlying manifold.
Later, in Section 5, we will show that for a torus $B$, these tropical fluxes between tropical cycles are related to symplectic fluxes between their Lagrangian lifts in the symplectic manifold $X(B):=T^{*}B/T^{*}_{\mathbb{Z}}B$ in the sense of Section 3.
2.2.1. Tropical (co)homology
Tropical (co)homology was originally introduced in [tropical_homology_ikmz], with the idea of encoding information about the (co)homology of an algebraic variety in the tropical (co)homology of its tropicalization. This is done by enriching singular (co)chains with coefficients in the exterior powers of the integral lattices $T_{\mathbb{Z}}^{(*)}B$ of a tropical manifold $B$, as a way of recording data from the tropical structure. In our case $B$ is a tropical torus, and we have seen that this means the local systems of lattices $T_{\mathbb{Z}}^{(*)}B$ are constant, therefore these are global coefficients.
2.2.2. Tropical chains from algebraic equivalences
This section does not contain any original ideas, but extends constructions of Zharkov [zharkov_tropical_ceresa, Section 3.1] from the case of $1$-cycles to that of arbitrary $k$-cycles. We borrow some of the notation and terminology for the sake of consistency. Note that here $B$ could be any tropical manifold, not just a tropical torus.
Our first aim is to translate algebraic equivalence of cycles into the language of tropical (co)chains, in order to obtain tropical (co)homological obstructions to two cycles being algebraic equivalent. This requires two steps. The first is to associate to a tropical cycle a tautological tropical cycle, which is a cycle in the tropical cohomology (i.e., it has coefficients in exterior powers of $T^{*}_{\mathbb{Z}}B$). The second is, given an algebraic equivalence between two cycles, to construct a tropical chain realising the tropical homological equivalence between the corresponding tautological tropical cycles.
Definition 2.8.
A tropical $k$-cycle in $B$ is a weighted balanced $k$-dimensional polyhedral complex with $\Gamma_{2}$-rational slopes.
Two such cycles $Z_{1}$ and $Z_{2}$ are algebraically equivalent in $B$ if there exists two points $c_{1}$ and $c_{2}$ in some tropical curve $C$, and a tropical $(k+1)$-cycle $W$ in $B\times C$ satisfying
To any tropical $k$-cycle $Y$, one can associate a canonical element of $C_{k}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ called a tautological tropical cycle $\overline{Y}$ as follows. Every oriented $k$-face $f$ of $Y$ with weight $w_{f}$ defines a $k$-cell $w_{f}\sigma_{f}$, as well as $k$ primitive tangent vectors $v_{1},\dots,v_{k}$ up to a transformation in $SL(n;\mathbb{Z})$. The data of an orientation on $f$ yields a preferred element $\beta(f)=v_{i_{1}}\wedge\dots\wedge v_{i_{k}}$ in $\Lambda^{k}T_{\mathbb{Z}}B$. Then the tautological cycle associated to $Y$ is
$\overline{Y}=\sum_{f\in Y}\beta(f)w_{f}\cdot\sigma_{f}.$
Reversing the orientation on $f$ has the effect of taking $\beta(f)$ to $\beta(f)$ as well as $\sigma_{f}$ to $\sigma_{f}$, and these cancel out in $C_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)$, making $\overline{Y}$ well-defined. The fact that $\overline{Y}$ is closed follows from the fact that $Y$ satisfies a balancing condition: an (n1)(n-1)-face in the boundary of $s$ distinct $n$-faces $f_{1},\dots,f_{s}$ will appear $s$ times with coefficients $\beta(f_{1})\omega_{f_{1}},\dots,\beta(f_{s})\omega_{f_{s}}$, whose sum vanishes by the balancing condition.
The existence of an algebraic equivalence $W\subset B\times C$ between two $k$-cycles $Y_{1}$ and $Y_{2}$ in $C$ implies that $\overline{Y_{1}}]=[\overline{Y_{2}}]$ in $H_{k}(B;\Lambda^{k}T_{\mathbb{Z}}B)$. Still following [zharkov_tropical_ceresa, Section 3.1], we construct some $\tilde{W}$ in $C_{k+1}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ satisfying
$\partial\tilde{W}=\overline{Y_{1}}-\overline{Y_{2}}.$
We proceed by choosing a path $P$ in $C$ from $c_{1}$ to $c_{2}$, and letting
$Y_{c}:=\pi_{*}(W\cap(B\times\{c\}))$
for any $c\in C$. Now define the weighted polyhedral complex $W_{P}\subset B\times C$ as the restriction of $W$ to $(\pi^{\prime})^{-1}(P)$, where $\pi^{\prime}:B\times C\longrightarrow C$ is the projection onto the second factor. For any $(k+1)$-cell $\sigma$ in the support of $W_{P}$, two scenarios arise. Either $\pi^{\prime}$ is not submersive on $\sigma$, in which case we set $\beta(\sigma)=0$, or it is submersive on $\sigma$. In this latter case, above each $c\in\pi^{\prime}(\sigma)$, $\sigma\cap(Y_{c}\times\{c\})$ carries a tautological framing $\beta_{c}(\sigma)\in\Lambda^{k}T_{\mathbb{Z}}(B\times\{c\})$ as decribed above, which can be pulled back by $\pi$ to $\Lambda^{k}T_{\mathbb{Z}}(B\times C)$. Clearly, on a fixed $(k+1)$-cell $\sigma$, this is independent of $c$ because both $\pi$ and $\pi^{\prime}$ are linear; denote this by $\beta(\sigma)$. We can now define
Lemma 2.9.
$\partial\tilde{W}=\overline{Y_{1}}-\overline{Y_{2}}.$
2.2.3. Determinantal forms
The notion of a determinantal form on a tropical torus $B$ was originally introduced by Zharkov [zharkov_tropical_ceresa, Section 3.2] as a cocycle in $C^{2}(B;T^{*}_{\mathbb{Z}}B)$. We extend this construction to higher dimensions. Although in practice we will only be interested in determinantal forms in $C^{k+1}(B;\Lambda^{k}T^{*}B)$, we construct them in any $C^{j}(B;\Lambda^{k}T^{*}B)$. In theory, one could expect a version of these determinantal forms to make sense in any tropical manifold, however our construction is for tropical tori.
Definition 2.10.
[…]
Remark 2.11.
$\Omega^{3}_{1,2}=e_{1}^{*}\otimes dx_{2}\wedge dx_{3}-e_{2}^{*}\otimes dx_{1}\wedge dx_{3}+e_{3}^{*}\otimes dx_{1}\wedge dx_{2}.$
The period lattices of the forms $\Omega^{n}_{k,j}$ are nicely packaged in the matrix $Q\in GL(n;\mathbb{R})$ determining the tropical torus $B=B(Q)$. We denote them by $\mathcal{P}^{n}_{k,j}\subset\mathbb{R}$.
Proposition 2.12.
The periods $\mathcal{P}^{n}_{j,k}$of the determinantal form $\Omega^{n}_{j,k}$ on a tropical torus $B(Q)$ are given by integer combinations of the real numbers
[…]
for any tuples $(i_{1}<\dots<i_{k})\in\{1,\dots,n\}^{k}$ and $(r_{1}<\dots<r_{j})\in\{1,\dots,n\}^{j}$, where $\sigma$ is the permutation taking $(i_{1},\dots,i_{k},s_{1},\dots,s_{j})$ to its corresponding ordered $(k+j)$-tuple, and $Q_{(s_{1},\dots,s_{j}),(r_{1},\dots,r_{j})}$ is the $j\times j$ minor of the $n\times n$ matrix $Q$ obtained by conserving only the rows $\{s_{1},\dots,s_{j}\}$ and the columns $\{r_{1},\dots,r_{j}\}$.
Proof.
The pairing of each term of the form
$dx_{s_{1}}\wedge\dots\wedge dx_{s_{j}}(\gamma_{r_{1}}\wedge\dots\wedge\gamma_{r_{j}})=dx_{s_{1}}\wedge\dots\wedge dx_{s_{j}}((Qe_{r_{1}})\wedge\dots\wedge(Qe_{r_{j}})).$
This is given by
$Q_{(s_{1},\dots,s_{j}),(r_{1},\dots,r_{j})}.$
Therefore the periods of the determinantal forms are generated by appropriate linear combinations of the minors of $Q$. ∎
Example 2.13.
In the case of the non-hyperelliptic genus $3$ curve from Section 2.1.3, these periods of $\Omega^{3}_{2,1}$ were computed by Zharkov [zharkov_tropical_ceresa, Section 3.2]. We use the bases represented in Figure 6. In particular $H_{2}(J(C);T_{\mathbb{Z}}J(C))\cong\Gamma_{2}\otimes\Lambda^{2}\Gamma_{1}$ has $9$ generators
$e_{1}\otimes\gamma_{1}\wedge\gamma_{2}$ $e_{2}\otimes\gamma_{1}\wedge\gamma_{2}$ $e_{3}\otimes\gamma_{1}\wedge\gamma_{2}$
$e_{1}\otimes\gamma_{1}\wedge\gamma_{3}$ $e_{2}\otimes\gamma_{1}\wedge\gamma_{3}$ $e_{3}\otimes\gamma_{1}\wedge\gamma_{3}$
$e_{1}\otimes\gamma_{2}\wedge\gamma_{3}$ $e_{2}\otimes\gamma_{2}\wedge\gamma_{3}$ $e_{3}\otimes\gamma_{2}\wedge\gamma_{3},$
and the matrix $Q$ is given by
$Q=\begin{pmatrix}a+e+f&-f&-e\\ -f&b+d+f&-d\\ -e&-d&c+d+e\end{pmatrix}$.
Using Proposition 2.12, these yield $6$ different periods, corresponding to the different minors of $Q$:
$ab+ad+af+be+de+ef+bf+df$ $ad+ed+df+ef$
$ac+ad+ae+ec+ed+cf+df+ef$ $df+be+de+ef$
$bc+bd+be+cd+de+cf+df+ef$ $cf+df+ef+de.$
This coincides with those obtained by Zharkov.
Remark 2.14.
$\phi_{Q}:\hskip 2.84526pt$ $\mathbb{Z}^{n}\longrightarrow\mathbb{R}^{n}$
$x\mapsto Q\cdot x$
viewed as a map from $H_{1}(B;\mathbb{Z})$ to $\Gamma_{1}\otimes\mathbb{R}\cong H^{0}(T_{\mathbb{R}}B(Q))$ is precisely the integration pairing, that is for any $\eta\in H^{0}(T_{\mathbb{R}}^{*}B(Q))$ and $\gamma\in H_{1}(B(Q);\mathbb{Z})$,
$\int_{\gamma}\eta=\eta(\phi_{Q}(\gamma)).$
Integration over $H_{k}(B(Q);\mathbb{Z})$ is obtained from exterior powers $\phi_{Q}$. In the case of a Jacobian $B=J(C)$, the image of this map is contained in the lattice $H^{0}(T_{\mathbb{Z}}J(C))\subset H^{0}(T_{\mathbb{R}}J(C))$. Very generally, as described in [nick_ivan1, Lemma 3.9], any polarisation
$c:H_{1}(B;\mathbb{Z})\longrightarrow H^{0}(T_{\mathbb{Z}}^{*}B)$
on a tropical torus can be related to its integration pairing
$\phi:H_{1}(B;\mathbb{Z})\longrightarrow H^{0}(T_{\mathbb{R}}B)$
through an isomorphism $T_{\mathbb{R}}B\cong T_{\mathbb{R}}^{*}B$ whose image lives in $T_{\mathbb{Z}}^{*}B(Q)$. In the case of the canonical principal polarisation on a Jacobian $J(C)$, one can verify that this isomorphism is simply the canonical duality isomorphism $T_{\mathbb{Z}}J(C)\cong T_{\mathbb{Z}}^{*}J(C)$.
2.2.4. Tropical flux
Recall that the $k$-th tropical Griffiths group $\text{Griff}_{k}(B)$ is defined as the set of homologically trivial tropical $k$-cycles in $B$, modulo the equivalence relation generated by algebraic equivalence.
If $dim(B)=n$, we define for all $k\geq 1$ the tropical flux maps
$\Theta_{k}:\hskip 2.84526pt\text{Griff}$ $_{k}(B)\longrightarrow\mathbb{R}/\mathcal{P}^{n}_{k+1,k}$
$\hskip 2.84526ptZ\hskip 14.22636pt\mapsto\hskip 5.69054pt\int_{\gamma}\Omega^{n}_{k+1,k},$
where $\gamma\in C_{k+1}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ satisfies $\partial\gamma=\overline{Z}$, and $\mathcal{P}^{n}_{k+1,k}$ is the period lattice of $\Omega^{n}_{k,k+1}$ as computed in Proposition 2.12.
The following Proposition and its subsequent Corollary 2.16 constitute the proof that this map is well-defined: while it is clear that a different choice of $(k+1)$-chain $\gamma$ would only change $\int_{\gamma}\Omega^{n}_{k+1,k}$ by a period of $\Omega^{n}_{k+1,k}$, it remains to show that for $Z$ algebraically trivial, this integral vanishes.
Proposition 2.15.
Given any integers $j$, $k$, and $n$ with $j+k\leq n$, a point $b$ in $B$, and vectors $v_{1},\dots,v_{j+k}$ in $T_{b}B$, we have
$(\Omega^{n}_{j,k})_{b}(v_{1},\dots,v_{j+k})=0$
Proof.
Let $J$ denote the vector space generated by the last $j$ vectors. Assume one of the last $j$ vectors $v_{k+l}$ is a linear combination of the other j+k1j+k-1 vectors. By the antisymmetry properties of $\Omega^{n}_{j,k}$,
$(\Omega^{n}_{j,k})_{b}(v_{1},\dots,v_{j+l},\dots,v_{j+k})=(\Omega^{n}_{j,k})_{b}(v_{1},\dots,v_{j+l}-\pi_{J}(v_{j+l}),\dots,v_{j+k}).$
This means that, without loss of generality, we can assume $v_{k+l}$ is written as a linear combination of the first $k$ vectors:
$v_{k+l}=\sum_{r=1}^{k}a_{r}v_{r}$
for some real numbers $a_{1},\dots,a_{k}$. For the same reason, by replacing say $v_{1}$ by $\frac{1}{a_{1}}v_{k+l}$, one can assume $v_{k+l}=a_{1}v_{1}$ is a multiple of $v_{1}$. To show the vanishing of $(\Omega^{n}_{j,k})_{b}(v_{1},\dots,v_{j+k})$, one writes down explicitly
$(\Omega^{n}_{j,k})_{b}(v_{1},\dots,v_{j+k})=\sum_{(i_{1}<\dots<i_{j+k})\in\{1,\dots,n\}^{j+k}\atop(j_{1}<\dots<j_{k})\in\{i_{1},\dots,i_{j+k}\}^{k}}$ $(-1)^{sgn(\sigma)}\left(\sum_{\mu\in S_{k}}(-1)^{sgn(\mu)}v_{1,\mu(j_{1})}\dots v_{k,\mu(j_{k})}\right)\times$
$\left(\sum_{\nu\in S_{j}}(-1)^{sgn(\nu)}v_{k+1,\nu(s_{1})}\dots v_{k+j,\nu(s_{j})}\right).$
Recall that $\sigma$ appears in the definition of $\Omega^{n}_{j,k}$ and is the permutation taking $\{i_{1},\dots,i_{j+k}\}$ to $\{j_{1},\dots,j_{k},s_{1},\dots,s_{j}\}$. When replacing $v_{k+l}$ by $a_{1}v_{1}$, one can see that for every term, composing $\sigma$ with the transposition $(j_{1}s_{l})$ yields a term equal in absolute value but with opposite sign, hence they cancel out in pairs. The same argument can be carried out by taking one of the first $k$ and assuming it is written as a linear combination of the last $j$ vectors. ∎
$(\Omega^{n}_{j,k})_{b}(v_{1},\dots,v_{n})=vol(v_{1},\dots,v_{n}).$
Corollary 2.16.
Assume $Z$ is an algebraically trivial $k$-cycle for $k\geq 1$. Then
$\int_{\gamma}\Omega^{n}_{k+1,k}\in\mathcal{P}^{n}_{k+1,k}$
Proof.
Let $W$ be an algebraic equivalence of $k$-cycles $Z_{1}$ and $Z_{2}$ for $k\geq 1$, and $\tilde{W}$ its associated $(k+1)$-chain in $C_{k+1}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ constructed in Section 2.2.2. Then it suffices to show
$\int_{\tilde{W}}\Omega^{n}_{k+1,k}=0.$
To compute this integral we locally pair $\Omega^{n}_{j,k}$ on each $k+1$-cell $\sigma$ in the support of $\tilde{W}$ with $k$ vectors coming from the framing $\beta(\sigma)$ and $k+1$ vectors tangent to $\sigma$. Notice that by construction the framing $\beta(\sigma)\in\Lambda^{k}T_{\mathbb{Z}}B$ is the pullback of a tautological framing on a $k$-cell. Of course a tautological framing is written as a wedge product of vectors living inside the tangent space to the $k$-cell, therefore they are clearly not linearly independent, and the integral vanishes. Notice that, if $k=0$, the $\beta(\sigma)$ are simply integers hence the linear dependence argument cannot be carried out. ∎
Remark 2.17.
We have actually proven a statement stronger than
$\int_{\tilde{W}}\Omega^{n}_{k+1,k}=0,$
namely that $\Omega^{n}_{k+1,k}|_{\tilde{W}}=0$, in the sense that $\Omega^{n}_{k+1,k}$ evaluates to zero at each point of $\tilde{W}$.
Remark 2.18.
We have mentioned that this result does not hold for $k=0$; the argument of linear dependence used for the vanishing result (Corollary 2.16) cannot be used for points. As a counterexample, consider the case $n=1$ where $B\cong S^{1}$ with two points $p$ and $q$ as shown below. In this setting $\Omega^{1}_{1,0}=dx\in C^{0}(B;T_{\mathbb{Z}}^{*}B)$.
$\int_{\tilde{\Delta}}\Omega^{1}_{1,0}=l(E)\neq 0.$